The difference between two numbers is 40. Four times the smaller is equal to 5 more than the larger. What are the numbers?

Answer by fractalier(6550)

Call the larger and smaller numbers x and y.
Then we have
x - y = 40 or x = y + 40
and
4y = x + 5
Substitute the first into the second and get
4y = y + 40 + 5
3y = 45
y = 15 and then
x = 55